OpenStudy (anonymous):

Beginning with the function f(x)=logax , state which transformations were used to obtain the functions given below: g(x)=loga(3−x)+1 t(x)=2loga3x The "a" is small

OpenStudy (anonymous):

Is that \[\log_a\]or \[\log(a...)\]

OpenStudy (australopithecus):

the a is small meaning it is subscript jabber

OpenStudy (anonymous):

if a is base. \[g \left( x \right)=\log_{a} \left( 3-x \right)+1=\log_{a} \left( 3-x \right)+\log_{a}a= \log_{a}a \left( 3-x \right) \] so if we replace x by a(3-x) we get the reqd.result.